What is the acceleration of a car (km/hr2) that goes from 40 km/hr to 80 km/hr in 0.3 hrs


1
Expert's answer
2021-01-15T13:46:37-0500

By the definition of acceleration, we have:


"a=\dfrac{v-v_0}{t},""a=\dfrac{80\ \dfrac{km}{hr}-40\ \dfrac{km}{hr}}{0.3\ hr}=133.3\ \dfrac{km}{hr^2}."

Answer:

"a=133.3\ \dfrac{km}{hr^2}."